# Understanding Fine-tuning in Approximate Unlearning: A Theoretical Perspective

### v2

@@ 3.2 Overlapping Features
Similarly to 1, $d_{r}$ and $d_{f}$ are the distinct feature numbers for remaining and forgetting data, while the equation $d_{r}+d_{lap}+d_{f}=d$ holds. It can also indicates that the optimal solution can be decomposed into $\mathbf{w}_{*}=\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap}+\mathbf{w}_{*}^{r}$ such that $\mathbf{y}^{f}=\mathbf{X}_{f}^{\top}(\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap})$ and $\mathbf{y}^{r}=\mathbf{X}_{r}^{\top}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$.
Theorem 2.
Suppose a model is trained by the procedure 2 and 3 separately. Under the 2, it holds that
@@
Theorem 2 shows that the previous conclusions remain valid under the assumptions of overlapping features, as the information from all training data, including forget data, is preserved from the pretrained model, $\mathbf{w}_{o}$, to the unlearned model through fine-tuning, $\mathbf{w}_{t}$. Consequently, the loss on both the remaining dataset and the forgetting dataset for the fine-tuning model is zero. Additionally, an interesting observation is that the number of overlapping features does not impact the unlearning accuracy of the fine-tuning model. The proof of Theorem 2 is provided in Section B.3.
Both Theorem 1 and Theorem 2 present similar findings regarding the performance of the unlearned model through fine-tuning. We run a synthetic experiment to validate these results (more experimental details in Appendix A). In Figure 1(a) and Figure 1(b), both distinct and overlapping feature assumptions demonstrate the same results: 1) The remaining loss of fine-tuning model $\mathbf{w}_{t}$ and golden model $\mathbf{w}_{g}$ is zero, indicating that the fine-tuning model performs equivalently to the golden model, successfully retaining the model’s utility on the remaining dataset. 2) The unlearning loss of the fine-tuning model consistently remains at zero, differing from the golden model, suggesting that the fine-tuning model fails to forget the information obtained from the pretrained model. These empirical findings align well with our theoretical analysis.
@@ 4 Eliminating Forgetting Data Features from Pre-Trained Model Enhances Unlearning
To mitigate the UA gap between the fine-tuning model and the golden model, it becomes evident that the remaining portion of the pretrained model does not contribute to UA. Specifically, the components of the pretrained model $\mathbf{w}_{o}$ associated with the forgetting data $(\mathbf{w}_{*}^{f})$ do not enhance performance on the remaining dataset $D_{r}$. Therefore, if we can eliminate the forgetting component—specifically by removing the $\mathbf{w}_{*}^{f}$ term from $\mathbf{w}_{o}$—the divergence can be addressed. In the following, we provide a formal description of this modification. Consider the same learning procedure Equation 1 to obtain the pretrained model $\mathbf{w}_{o}$. Prior to unlearning through fine-tuning, we modify $\mathbf{w}_{o}$ by removing components associated with the forgetting data. Specifically, we construct a modified model $\hat{\mathbf{w}}_{o}$ as follows:
@@
    In the distinct features scenario (i.e., under the 1), we have:
@@
    In the overlapping features scenario (i.e., under 2), we have:
@@
According to Theorem 3, under the distinct features assumption, the masked unlearned model achieves the same remaining and unlearning loss as the golden model. Furthermore, when considering overlapping features, if the overlapped component from the pretrained model is retained, the remaining loss remains zero, as with the golden model, while the unlearning loss differs only in the projection component. This difference can be considered negligible when applied to $\mathbf{w}_{*}^{r}$ and $\mathbf{w}_{*}^{lap}$ due to the model assumption. Figure 1(c) and Figure 1(d) verify our theoretical conclusions. However, if the overlapped component is discarded from the pretrained model, the remaining loss is no longer zero, and there is a small change to the unlearning loss that can be overlooked. These findings offer several insights into the design of machine unlearning algorithms: 1) Masking on the pretrained model can significantly improve unlearning accuracy while preserving the retaining accuracy. If we can identify the component of the pretrained model related to the forgetting data, applying masking to this component can further enhance UA. Our theorem also explains recent related works, such as [17, 6], which apply a mask to the pretrained model either randomly or by regularizing the weights associated with the forgetting data to provide better unlearn performance. These methods share the same underlying principle discussed here. However, one overlook a critical scenario: when the remaining data and forgetting data share similar features, it becomes unclear whether those shared features should be retained. As shown in Theorem 3, our work provides a clear resolution to this question. 2) When considering overlapping features, retaining them does not substantially affect unlearning accuracy, but discarding them compromises the retaining accuracy. As shown in Theorem 3, the remaining loss can not retain zero unless the remaining data $\mathbf{X}_{r}$ can be fully represented by the fine-tuning space, meaning $\mathbf{P}_{t}\mathbf{X}_{r}=\mathbf{X}_{r}$. Additionally, as the number of overlapping features increases, the impact on both remaining and unlearning loss becomes more significant. Discarding too many overlapping components can lead to instability in the retaining accuracy, as the model loses essential information needed to represent $D_{r}$, which in turn causes the remaining loss to increase. Figure 1 and Figure 2 validate our theoretical findings. The proof of Theorem 3 is provided in Section B.4.
Figure 2: Comparison of Machine Unlearning Loss with and without Overlapping Features. Figure 2(a) retains overlapping features from the pretrained model, showing the matching performance between masked $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; Figure 2(b) discards the overlapping features, showing a decline in retaining accuracy.
@@ 5.1 Discarding Overlapping Features May Harm Retaining Accuracy
In Section 4 we show that once the components of the pretrained model related to the forgetting data are identified and removed, unlearning accuracy can be significantly improved. In practice, an existing work [6] shares the same principles as discussed above. Specifically, it constructs the desired weight saliency map by leveraging the gradient of a loss function with respect to the model weights on the forgetting dataset:
@@
where $\mathbbm{1}(g\geq\gamma)$ is an element-wise indicator function that outputs 1 for the $i$-th element if $g_{i}\geq\gamma$ and 00 otherwise, ||$\cdot$|| ⋅ | denotes the element-wise absolute value operation, and $\gamma>0$ is a hard threshold. Variables with larger gradients, which are highly associated with the forgetting dataset, are identified and subsequently masked. Then, the unlearn model can be updated by:
$\mathbf{w}_{u_{f}}=\underbrace{\mathbf{m}_{f}\odot(\Delta\mathbf{w}+\mathbf{w}_{o})}_{\text{salient weights }}+\underbrace{(\mathbf{1}-\mathbf{m}_{f})\odot\mathbf{w}_{o}}_{\text{original weights }},$ (7)
@@
However, Equation 7 overlooks the scenario where the remaining data and the forgetting data share similar model weights, making it unclear whether those weights should be preserved. In Theorem 3, we demonstrate that when considering overlapping features (features shared between the remaining dataset and the forgetting dataset), retaining them is more beneficial than discarding them, which contrasts slightly with the approach in [6]. Specifically, we show that discarding overlapping features compromises retaining accuracy while retaining them ensures better performance for the remaining dataset.
@@
$\mathbf{w}_{u_{r}}=\underbrace{(\mathbf{1}-\mathbf{m}_{r})\odot(\Delta\mathbf{w}+\mathbf{w}_{o})}_{\text{salient weights }}+\underbrace{\mathbf{m}_{r}\odot\mathbf{w}_{o}}_{\text{original weights }}.$ (10)
Here, $\Delta\mathbf{w}$ is optimized using Equation 8 with the retention-based mask, which combines the forgetting loss on the random labels of forgotten data with the retaining loss on the remaining data.
@@ 5.2 Experiment
Datasets and Models. Our baseline methods include the naive fine-tuning approach [8, 29] and SalUn [6], both implemented on ResNet-18 [12]. For comparison, we also include the golden retrained model as a reference. Our experiments will focus on image classification using the CIFAR-10 [15] and CIFAR-100 [15]. More details on the experimental setup will be provided in Appendix A.
Evaluation Metrics. We follow the existing work to assess machine unlearning performance from different aspects [8, 10, 27, 17, 23, 32]. Specifically, we focus on the following evaluation metrics:
  • Unlearning accuracy (UA): We define $\operatorname{UA}(\mathbf{w}_{t})=1-\operatorname{Acc}_{D_{f}}(\mathbf{w}_{t})$ as [17], measuring how effectively the model has forgotten the targeted data. Here $\operatorname{Acc}_{D_{f}}(\mathbf{w}_{t})$ is the accuracy of the unlearned model on the forgetting dataset.
@@
Retention-Based Masking Improves Retaining Accuracy. Next, we compare retention-based masking (Ours) and forgetting-based masking (SalUn) to evaluate how different masking approaches impact retaining accuracy, measured by the distance from the retrained model. For CIFAR-10 Random Forgetting, SalUn achieves a gap of $0.38$, while RBM improves it to $0.28$. For CIFAR-100 Random Forgetting, SalUn shows a gap of $1.33$, whereas RBM significantly reduces it to $1.24$. In class-wise forgetting scenarios, RBM consistently maintains an advantage: for CIFAR-10 Classwise Forgetting, SalUn achieves a gap of $0.53$ compared to $0.31$ for RBM. Similarly, for CIFAR-100 Class-wise Forgetting, SalUn achieves a gap of $0.35$, while RBM further improves this to $0.08$. These results underscore that the retention-based saliency approach is more effective in preserving the performance of the remaining dataset, aligning with our theoretical insights in Theorem 3.
@@
6 Conclusion
In conclusion, we present the first theoretical analysis of fine-tuning methods for machine unlearning within a linear regression framework. Our analysis, covering two scenarios—distinct and overlapping feature sets—demonstrates that while fine-tuning can achieve optimal retaining accuracy, it fails to fully unlearn the forgetting dataset. This failure arises from the pretrained model retaining information about the forgetting data. To address this, we propose a theoretical solution and introduce Retention-Based Masking (RBM), a strategy that constructs masks based on the remaining dataset to preserve overlapping features. Our experimental results demonstrate that RBM achieves a better balance between unlearning and retaining objectives compared to existing masking approaches.
Appendix A Experimental Details
A.1 Verification via Simulation
To empirically validate the theoretical findings presented in Theorem 1, Theorem 2 and Theorem 3 regarding the performance of unlearned models through fine-tuning, we first conducted a series of synthetic experiments.
Data Generation. We constructed two data matrices, $\mathbf{X}_{r}$ and $\mathbf{X}_{f}$, representing the remaining data and the forgetting data, respectively. The remaining data matrix, $\mathbf{X}_{r}^{\top}$, was structured as $[\mathbf{R}^{\top},\mathbf{L}_{1}^{\top},\mathbf{0}]$, and the forgetting data matrix, $\mathbf{X}_{f}^{\top}$, as $[\mathbf{0},\mathbf{L}_{2}^{\top},\mathbf{F}^{\top}]$, where $\mathbf{L}_{1}^{\top}=\mathbf{L}_{2}^{\top}=\mathbf{0}$ to enforce the non-overlapping case. Here, $\mathbf{R}^{\top}$ and $\mathbf{F}^{\top}$ are random matrices corresponding to different feature sets, and the zeros represent the distinct features across the datasets. We set the total number of data points to $n=40$ and the total number of features to $d=40$. The remaining data consisted of $n_{r}=30$ samples with $d_{r}=20$ features, while the forgetting data comprised $n_{f}=10$ samples with $d_{f}=d-d_{r}=20$ features. To simulate a controlled environment, we fixed the number of overlapping features to $d_{lap}=0$ and $d_{lap}=8$ for non-overlapping case and overlapping case, respectively.
Label Generation. We generated the true coefficient vector $\mathbf{w}_{*}\in\mathbb{R}^{d}$ by sampling from a standard normal distribution. The labels were created using a linear regression model without added noise:$\mathbf{y}=\mathbf{X}^{\top}\mathbf{w}_{*}$. The labels were partitioned into $\mathbf{y}_{r}$ and $\mathbf{y}_{f}$, corresponding to the remaining and forgetting data, respectively.
Model Training. To compare the effects of fine-tuning, we considered two models: the fine-tuning model $\mathbf{w}_{t}$ and the golden model $\mathbf{w}_{g}$. Specifically, $\mathbf{w}_{t}$ was obtained by fine-tuning on a subset of the remaining data, denoted as $\mathbf{X}_{t}$, which consisted of the first $n_{t}$ data points from $\mathbf{X}_{r}$. The value of $n_{t}$ varied from 1 to $n_{r}-1$ to study the impact of the fine-tuning data size. The initial model $\mathbf{w}_{o}$ was derived from the entire dataset $\mathbf{X}$ and calculated by the Equation 1. $\mathbf{w}_{g}$ was trained from scratch on the entire remaining data $\mathbf{X}_{r}$ and computed by solving Equation 2. If considering the masked case in synthetic data, the masked pretrained model will be constructed by zeroing out the coefficients corresponding to the forgetting data features with (without) overlapping features.
  • Remaining Data Loss (RA Loss):$\operatorname{MSE}_{\mathrm{RA}}(\mathbf{w})=\frac{1}{n_{r}}\|\mathbf{X}_{r}\mathbf{w}-\mathbf{y}_{r}\|^{2}$
  • Unlearning Data Loss (UA Loss):$\operatorname{MSE}_{\mathrm{UA}}(\mathbf{w})=\frac{1}{n_{f}}\|\mathbf{X}_{f}\mathbf{w}-\mathbf{y}_{f}\|^{2}$.
Experimental Results Figure 1(c) and Figure 1(d) illustrate that the masked fine-tuning method discussed in Section 4 can significantly improve unlearning accuracy while preserving the retaining accuracy. Specifically, both the remaining loss and unlearning loss of $\hat{\mathbf{w}}_{t}$ perfectly match those of the golden model under both distinct and overlapping feature scenarios. Additionally, Figure 2 present comparisons of machine unlearning loss for different approaches to handling overlapping features: Figure 2(a) retains overlapping features from the pretrained model, demonstrating matching performance between the masked $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; whereas Figure 2(b) discards the overlapping features, resulting in a decline in retaining accuracy. These empirical results align well with our theoretical findings.
A.2 Additional Real-world Details
@@
Visualization of Remaining Data and Forgetting Data Features. The visualization in Figure 3 uses t-SNE to project feature representations of the forgetting and remaining data in CIFAR-10 and CIFAR-100 datasets. The red points correspond to forgetting data, and the blue points represent remaining data. This visualization aims to demonstrate that, in class-wise datasets, the unlearning task for a specific class may involve distinct features. In such cases, naive fine-tuning (FT) methods tend to contribute less towards forgetting the class and focus more on retaining features from the pretrained model.
Appendix B Proofs
B.1 Useful properties

### v3

@@ 3.2 Overlapping Features
Similarly to ˜3.1, $d_{r}$ and $d_{f}$ are the distinct feature numbers for remaining and forgetting data, while the equation $d_{r}+d_{lap}+d_{f}=d$ holds. It can also indicates that the optimal solution can be decomposed into $\mathbf{w}_{*}=\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap}+\mathbf{w}_{*}^{r}$ such that $\mathbf{y}^{f}=\mathbf{X}_{f}^{\top}(\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap})$ and $\mathbf{y}^{r}=\mathbf{X}_{r}^{\top}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$.
Theorem 3.4.
Suppose a model is trained by the procedure 2 and 3 separately. Under the ˜3.3, it holds that
@@
Theorem˜3.4 shows that the previous conclusions remain valid under the assumptions of overlapping features, as the information from all training data, including forget data, is preserved from the pretrained model, $\mathbf{w}_{o}$, to the unlearned model through fine-tuning, $\mathbf{w}_{t}$. Consequently, the loss on both the remaining dataset and the forgetting dataset for the fine-tuning model is zero. Additionally, an interesting observation is that the number of overlapping features does not impact the unlearning accuracy of the fine-tuning model. The proof of Theorem˜3.4 is provided in Section˜C.3.
Both Theorem˜3.2 and Theorem˜3.4 present similar findings regarding the performance of the unlearned model through fine-tuning. We run a synthetic experiment to validate these results (more experimental details in Appendix˜B). In Figure˜1(a) and Figure˜1(b), both distinct and overlapping feature assumptions demonstrate the same results: 1) The remaining loss of fine-tuning model $\mathbf{w}_{t}$ and golden model $\mathbf{w}_{g}$ is zero, indicating that the fine-tuning model performs equivalently to the golden model, successfully retaining the model’s utility on the remaining dataset. 2) The unlearning loss of the fine-tuning model consistently remains at zero, differing from the golden model, suggesting that the fine-tuning model fails to forget the information obtained from the pretrained model. These empirical findings align well with our theoretical analysis.
@@ 4 Eliminating Forgetting Data Features from Pre-Trained Model Enhances Unlearning
To mitigate the UA gap between the fine-tuning model and the golden model, it becomes evident that the remaining portion of the pretrained model does not contribute to UA. Specifically, the components of the pretrained model $\mathbf{w}_{o}$ associated with the forgetting data $(\mathbf{w}_{*}^{f})$ do not enhance performance on the remaining dataset $D_{r}$. Therefore, if we can eliminate the forgetting component—specifically by removing the $\mathbf{w}_{*}^{f}$ term from $\mathbf{w}_{o}$—the divergence can be addressed. In the following, we provide a formal description of this modification. Consider the same learning procedure Equation˜1 to obtain the pretrained model $\mathbf{w}_{o}$. Prior to unlearning through fine-tuning, we modify $\mathbf{w}_{o}$ by removing components associated with the forgetting data. Specifically, we construct a modified model $\hat{\mathbf{w}}_{o}$ as follows:
@@
    In the distinct features scenario (i.e., under the ˜3.1), we have:
@@
    In the overlapping features scenario (i.e., under ˜3.3), we have:
@@
According to Theorem˜4.1, under the distinct features assumption, the masked unlearned model achieves the same remaining and unlearning loss as the golden model. Furthermore, when considering overlapping features, if the overlapped component from the pretrained model is retained, the remaining loss remains zero, as with the golden model, while the unlearning loss differs only in the projection component. This difference can be considered negligible when applied to $\mathbf{w}_{*}^{r}$ and $\mathbf{w}_{*}^{lap}$ due to the model assumption. Figure˜1(c) and Figure˜1(d) verify our theoretical conclusions. However, if the overlapped component is discarded from the pretrained model, the remaining loss is no longer zero, and there is a small change to the unlearning loss that can be overlooked. These findings offer several insights into the design of machine unlearning algorithms: 1) Masking on the pretrained model can significantly improve unlearning accuracy while preserving the retaining accuracy. If we can identify the component of the pretrained model related to the forgetting data, applying masking to this component can further enhance UA. Our theorem also explains recent related works, such as liu2024model; fan2023salun, which apply a mask to the pretrained model either randomly or by regularizing the weights associated with the forgetting data to provide better unlearn performance. These methods share the same underlying principle discussed here. However, one overlook a critical scenario: when the remaining data and forgetting data share similar features, it becomes unclear whether those shared features should be retained. As shown in Theorem˜4.1, our work provides a clear resolution to this question. 2) When considering overlapping features, retaining them does not substantially affect unlearning accuracy, but discarding them compromises the retaining accuracy. As shown in Theorem˜4.1, the remaining loss can not retain zero unless the remaining data $\mathbf{X}_{r}$ can be fully represented by the fine-tuning space, meaning $\mathbf{P}_{t}\mathbf{X}_{r}=\mathbf{X}_{r}$. Additionally, as the number of overlapping features increases, the impact on both remaining and unlearning loss becomes more significant. Discarding too many overlapping components can lead to instability in the retaining accuracy, as the model loses essential information needed to represent $D_{r}$, which in turn causes the remaining loss to increase. Figure˜1 and Figure˜2 validate our theoretical findings. The proof of Theorem˜4.1 is provided in Section˜C.4.
Figure 2: Comparison of Machine Unlearning Loss with and without Overlapping Features. Figure˜2(a) retains overlapping features from the pretrained model, showing the matching performance between masked $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; Figure˜2(b) discards the overlapping features, showing a decline in retaining accuracy.
@@ 5.1 Discarding Overlapping Features May Harm Retaining Accuracy
In Section˜4 we show that once the components of the pretrained model related to the forgetting data are identified and removed, unlearning accuracy can be significantly improved. In practice, an existing work fan2023salun shares the same principles as discussed above. Specifically, it constructs the desired weight saliency map by leveraging the gradient of a loss function with respect to the model weights on the forgetting dataset:
@@
where $\mathbbm{1}(g\geq\gamma)$ is an element-wise indicator function that outputs 1 for the $i$-th element if $g_{i}\geq\gamma$ and 0 otherwise, $|\cdot|$ denotes the element-wise absolute value operation, and $\gamma>0$ is a hard threshold. Variables with larger gradients, which are highly associated with the forgetting dataset, are identified and subsequently masked. Then, the unlearn model can be updated by:
[…]
@@
However, Equation˜7 overlooks the scenario where the remaining data and the forgetting data share similar model weights, making it unclear whether those weights should be preserved. In Theorem˜4.1, we demonstrate that when considering overlapping features (features shared between the remaining dataset and the forgetting dataset), retaining them is more beneficial than discarding them, which contrasts slightly with the approach in fan2023salun. Specifically, we show that discarding overlapping features compromises retaining accuracy while retaining them ensures better performance for the remaining dataset.
@@
[…]
Here, $\Delta\mathbf{w}$ is optimized using Equation˜8 with the retention-based mask, which combines the forgetting loss on the random labels of forgotten data with the retaining loss on the remaining data.
@@ 5.2 Experiment
Datasets and Models. Our baseline methods include the naive fine-tuning approach (golatkar2020eternal; warnecke2021machine) and SalUn (fan2023salun), both implemented on ResNet-18 (he2016deep). For comparison, we also include the golden retrained model as a reference. Our experiments will focus on image classification using the CIFAR-10 (krizhevsky2009learning) and CIFAR-100 (krizhevsky2009learning). More details on the experimental setup will be provided in Appendix˜B.
Evaluation Metrics. We follow the existing work to assess machine unlearning performance from different aspects (golatkar2020eternal; graves2021amnesiac; thudi2022unrolling; liu2024model; sharma2024discriminative; zhu2024decoupling). Specifically, we focus on the following evaluation metrics:
  • Unlearning accuracy (UA): We define $\operatorname{UA}(\mathbf{w}_{t})=1-\operatorname{Acc}_{D_{f}}(\mathbf{w}_{t})$ as liu2024model, measuring how effectively the model has forgotten the targeted data. Here $\operatorname{Acc}_{D_{f}}(\mathbf{w}_{t})$ is the accuracy of the unlearned model on the forgetting dataset.
@@
Retention-Based Masking Improves Retaining Accuracy. Next, we compare retention-based masking (Ours) and forgetting-based masking (SalUn) to evaluate how different masking approaches impact retaining accuracy, measured by the distance from the retrained model. For CIFAR-10 Random Forgetting, SalUn achieves a gap of $0.38$, while RBM improves it to $0.28$. For CIFAR-100 Random Forgetting, SalUn shows a gap of $1.33$, whereas RBM significantly reduces it to $1.24$. In class-wise forgetting scenarios, RBM consistently maintains an advantage: for CIFAR-10 Classwise Forgetting, SalUn achieves a gap of $0.53$ compared to $0.31$ for RBM. Similarly, for CIFAR-100 Class-wise Forgetting, SalUn achieves a gap of $0.35$, while RBM further improves this to $0.08$. These results underscore that the retention-based saliency approach is more effective in preserving the performance of the remaining dataset, aligning with our theoretical insights in Theorem˜4.1.
@@
Visualization of Remaining Data and Forgetting Data Features. The visualization in Figure˜3 uses t-SNE to project feature representations of the forgetting and remaining data in CIFAR-10 and CIFAR-100 datasets. The red points correspond to forgetting data, and the blue points represent remaining data. This visualization aims to demonstrate that, in class-wise datasets, the unlearning task for a specific class may involve distinct features. In such cases, naive fine-tuning (FT) methods tend to contribute less towards forgetting the class and focus more on retaining features from the pretrained model.
6 Conclusion
In conclusion, we present the first theoretical analysis of fine-tuning methods for machine unlearning within a linear regression framework. Our analysis, covering two scenarios—distinct and overlapping feature sets—demonstrates that while fine-tuning can achieve optimal retaining accuracy, it fails to fully unlearn the forgetting dataset. This failure arises from the pretrained model retaining information about the forgetting data. To address this, we propose a theoretical solution and introduce Retention-Based Masking (RBM), a strategy that constructs masks based on the remaining dataset to preserve overlapping features. Our experimental results demonstrate that RBM achieves a better balance between unlearning and retaining objectives compared to forgotten-based masking approaches.
Appendix A Discussion
Our analysis is limited to the overparameterized linear regression setting, which simplifies understanding but does not fully capture the complexity of modern nonlinear models. The assumptions of distinct and overlapping features are idealized and may not directly reflect real-world data distributions. Empirically, our evaluation is restricted to moderate-scale datasets and fine-tuning–based baselines, leaving broader comparisons to future work. Extending our framework to nonlinear or large-scale settings and exploring efficient or certified unlearning mechanisms are promising directions for future research.
Appendix B Experimental Details
B.1 Verification via Simulation
To empirically validate the theoretical findings presented in Theorem˜3.2, Theorem˜3.4 and Theorem˜4.1 regarding the performance of unlearned models through fine-tuning, we first conducted a series of synthetic experiments.
Data Generation. We constructed two data matrices, $\mathbf{X}_{r}$ and $\mathbf{X}_{f}$, representing the remaining data and the forgetting data, respectively. The remaining data matrix, $\mathbf{X}_{r}^{\top}$, was structured as $\mathbf{R}^{\top},\mathbf{L}_{1}^{\top},\mathbf{0}]$, and the forgetting data matrix, $\mathbf{X}_{f}^{\top}$, as $\mathbf{0},\mathbf{L}_{2}^{\top},\mathbf{F}^{\top}]$, where $\mathbf{L}_{1}^{\top}=\mathbf{L}_{2}^{\top}=\mathbf{0}$ to enforce the non-overlapping case. Here, $\mathbf{R}^{\top}$ and $\mathbf{F}^{\top}$ are random matrices corresponding to different feature sets, and the zeros represent the distinct features across the datasets. We set the total number of data points to $n=40$ and the total number of features to $d=40$. The remaining data consisted of $n_{r}=30$ samples with $d_{r}=20$ features, while the forgetting data comprised $n_{f}=10$ samples with $d_{f}=d-d_{r}=20$ features. To simulate a controlled environment, we fixed the number of overlapping features to $d_{lap}=0$ and $d_{lap}=8$ for non-overlapping case and overlapping case, respectively.
[…]
Model Training. To compare the effects of fine-tuning, we considered two models: the fine-tuning model $\mathbf{w}_{t}$ and the golden model $\mathbf{w}_{g}$. Specifically, $\mathbf{w}_{t}$ was obtained by fine-tuning on a subset of the remaining data, denoted as $\mathbf{X}_{t}$, which consisted of the first $n_{t}$ data points from $\mathbf{X}_{r}$. The value of $n_{t}$ varied from 1 to $n_{r}-1$ to study the impact of the fine-tuning data size. The initial model $\mathbf{w}_{o}$ was derived from the entire dataset $\mathbf{X}$ and calculated by the Equation˜1. $\mathbf{w}_{g}$ was trained from scratch on the entire remaining data $\mathbf{X}_{r}$ and computed by solving Equation˜2. If considering the masked case in synthetic data, the masked pretrained model will be constructed by zeroing out the coefficients corresponding to the forgetting data features with (without) overlapping features.
[…]
Experimental Results Figure˜1(c) and Figure˜1(d) illustrate that the masked fine-tuning method discussed in Section˜4 can significantly improve unlearning accuracy while preserving the retaining accuracy. Specifically, both the remaining loss and unlearning loss of $\hat{\mathbf{w}}_{t}$ perfectly match those of the golden model under both distinct and overlapping feature scenarios. Additionally, Figure˜2 present comparisons of machine unlearning loss for different approaches to handling overlapping features: Figure˜2(a) retains overlapping features from the pretrained model, demonstrating matching performance between the masked $\mathbf{w}_{t}$ model and golden model $\mathbf{w}_{g}$; whereas Figure˜2(b) discards the overlapping features, resulting in a decline in retaining accuracy. These empirical results align well with our theoretical findings.
B.2 Additional Real-world Details
Table 3: Unlearning Results on TinyImageNet. Performance comparison on the more challenging TinyImageNet dataset. Random 10% forgetting scenario with 50% masking. (RA: Retaining Accuracy, UA: Unlearning Accuracy, TA: Test Accuracy, MIA Efficacy: Membership Inference Attack Efficacy, Avg. Disparity: utility–privacy imbalance).
Method RA UA TA MIA Efficacy Avg. Disparity
Retrain 99.98 39.91 62.21 65.13
RBM 96.89 47.23 59.91 69.98 4.39
SalUn 97.20 31.30 59.10 71.08 5.11
We further evaluate our approach on the TinyImageNet dataset. This dataset is a subset of ImageNet, containing 200 classes with 500 training samples per class and 50 samples each for validation and testing. We consider a random forgetting scenario involving $10\%$ of the data, with a masking ratio of $50\%$ for both methods. The larger class space and higher visual diversity make this a particularly challenging benchmark. Nevertheless, our approach substantially outperforms the baseline method, showing only a moderate decline in test accuracy—reflecting the dataset’s inherent complexity. RBM maintains high retention accuracy while effectively mitigating MIA attacks to levels comparable to a retrained model, thereby achieving a lower average disparity relative to this standard.
Table 4: SVHN Unlearning Results using ViT Tiny. Evaluation of unlearning methods on SVHN. RA: Retaining Accuracy, UA: Unlearning Accuracy, TA: Test Accuracy, MIA Efficacy: Membership Inference Attack Efficacy.
Method RA UA TA MIA Efficacy Avg. Disparity
Retrain 100.00 9.76 89.38 15.27
RBM 94.12 8.15 88.73 13.82 2.39
SalUn 95.93 7.00 88.55 17.89 2.57
To further assess generality, we incorporate a new set of experiments using a Vision Transformer (ViT) backbone, extending beyond the conventional ResNet-based architectures commonly employed in prior work. Specifically, we adopt the ViT-Tiny model (approximately 5.7M parameters) and evaluate it on the SVHN dataset. We randomly forget $10\%$ of the data using a $50\%$ masking ratio. The learning rate for the ViT baseline is set to $1\times 10^{-3}$, and we perform a learning-rate ablation for SalUn by varying it between $6\times 10^{-4}$ and $2\times 10^{-3}$, obtaining the best performance at $1\times 10^{-3}$. For our method, the learning rate is adjusted to $5\times 10^{-4}$, yielding more stable convergence and consistent improvements across all metrics. As shown in Table 4, our proposed method consistently surpasses the forget-set–based unlearning method SalUn across all evaluated metrics.
Appendix C Proofs
C.1 Useful properties